HUBBLE SPACE TELESCOPE DAILY REPORT #2058 PERIOD COVERED: 0000Z (UTC) 02/10/98 - 0000Z (UTC) 02/11/98 Daily Status Report as of 042/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.3 Completed NIC/1/2/3 7831 (NICMOS Study of Nova Shells) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the nova shock shell around nova V1974-Cyg. The angular expansion rates and spatial distribution of the ejecta in both the continuum and emission lines of ionized gas will be determined. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Six Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.5 Completed STIS/CCD 7674 (IR Standards) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star HZ21 to establish a set of absolute flux standard on the WD scale that extend the existing wavelength coverage to 10300A with the STIS G750L grating setting. The proposal completed with no reported problems. 1.6 Completed WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.7 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.8 Completed Three Sets WF/PC-2 7505 (The Hubble Diagram for Distant Supernovae Measuring Cosmic Deceleration and Global Curvature) The WF/PC-2 was used to make observations of Super Nova segments SN1998D1, SN1998B1 and SN1998C1 in order to produce a precise Hubble diagram from measuring cosmic deceleration and global curvature. HST can make an essential contribution by providing photometry with substantially higher precision and better temporal coverage than can be obtained from the ground. The observations were completed as planned, and no anomalies were reported. 1.9 Completed STIS/CCD/WF/PC-2/NIC/1/2 7569 (The Nuclear Region of NGC-4151) The Space Telescope Imaging Spectrograph (CCD), the WF/PC-2 and the Near Infrared Camera (Cameras One and Two) were used to make observations of the nuclear region of Seyfert galaxy NGC-4151 which will be studied in several ways: 1) low-resolution spectra over the entire UV and visible range will be obtained; 2) slitless spectra of strong emission lines at moderate resolution and narrow-band WF/PC-2 images will be used to map the dynamics of the regions; 3) Echelle spectra over the entire UV bandpass will be used to study the intrinsic absorption in the nucleus; and 4) parallel imaging will be performed of the outer galaxy and other galaxies or QSOs. The observations were completed as planned, and no problems were reported. 1.10 Completed Three Sets of NIC/2 7327 (Elliptical Galaxy Evolution at High Redshift: IR Imaging of 60 Distant Clusters) The Near Infrared Camera (Camera Two) was used to make infrared observations high redshifted elliptical galaxies in galactic cluster CL+1058+463. By measuring the structural properties of high redshift cluster elliptical galaxies, it is hoped to resolve the controversy regarding the age and formation history of bright cluster ellipticals. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed FOC/96 7685 (FOC F/96 UV Throughput Monitoring) The Faint Object Camera (f/96) was used to make observations of standard star BPM16274 in order to perform ultraviolet throughput monitoring. These observations are taken four time during the cycle. An internal calibration flat image was also taken. The observations were completed as scheduled, and no problems were reported. 1.12 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 25 Successful : 25 2.3 Operations Notes: Flight operations personnel returned to the MOR from the SAMS backup facility at approximately 041/2330Z. No problems were encountered. As documented in HSTAR 6341, at 042/033807Z, the STIS MAMA1 MCE reset while in low voltage and while in the SAA. The error counter was reset. Science was not affected. MAMA1 recovery will be accomplished via the SMS. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH